OpenStudy (anonymous):

Christopher is 68 years old and Gabriela is 4 years old.How many years will it take until Christopher is only 5 times as old as Gabriela? (yes i know I have to set up an equation. but thats the tricky part)

OpenStudy (anonymous):

5(4+x).... that could be part of it :) do you see why?

OpenStudy (anonymous):

yeah but what about the other side of the equation?

OpenStudy (anonymous):

because i also have to include the 68 in there

OpenStudy (anonymous):

You can solve this question in two ways 1) Setting up an equation Christopher's age=68, let X no.of years pass before Christopher becomes 5 times as old as Gabriela..so Christopher's age after X years=68+X. Gabriela's age after X years=4+X. Now according to the criterion set, 68+X=5*(4+X) now solve for X.answer should be 12 2)Without setting up the equation Difference between Christopher's and Gabriela's Age=68-4=64 years Now. Christopher will at one point be only 5 time as old as Gabriela. At that point, Difference in ages= 5Gabriela-Gabriela=4Gabriela --> 4Gabriela=64 or Gabriela=16 years which means 16-4=12 more years need to pass

OpenStudy (anonymous):

okay thank you so much :]